OpenStudy (anonymous):

Mass of Vinegar 4.422g Volume of NaOH used 37.89 mL Moles of NaOH used 0.003789 C how do you find the: moles of HC2H3O2 neutralized, the mass of HC2H3O2, % HC2H3O2 in vinegar and the average % HC2H3O2 in vinegar?

OpenStudy (abmon98):

Construct a Balanced equation CH3COOH+NaOH-->CH3COONa+H2O Number of moles=mass/molar mass 4.422/(12*2)+(4*1)+(16*2)=0.0737 Ratio of moles are 1:1 so we use the limiting reagent which is Sodium hydroxide and Ethanoic acid(vinegar)is in excess. 0.0737-0.003789=0.069911 moles of CH3COOH were left unreacted. molesof CH3COOH neutralized were 0.003789

OpenStudy (abmon98):

0.069911*60=4.19466 grams did not react.

OpenStudy (abmon98):

(4.19466/4.422)*100=94.858%